# Latent Representation Matters: Human-like Sketches in One-shot Drawing Tasks

### v1

@@ 2 Related work
Human-machine comparison in one-shot drawing tasks: Cognitive scientists have developed various methods to compare the generalization abilities of machines and brains on drawing tasks. Lake et al. [44] introduced the Omniglot challenge in which both humans and machines are tasked with drawing symbols from categories represented by a single exemplar (see [45] for a review on the challenge). The authors evaluated the drawings’ recognizability in a visual Turing test where humans (or classifiers) had to distinguish between human-drawn and machine-generated symbols [11]. Additional metrics, including classification uncertainty and semantic similarity, were also used to compare drawings produced by humans and machines under different time constraints [46, 8]. While these evaluation frameworks provide useful insights into a sample’s recognizability, they do not measure how the diversity of model-generated samples compares to those created by humans. The “originality vs. recognizability” framework [31] mitigates this issue by adding the originality metric. An originality score quantifies the similarity between the original exemplar and its corresponding variations (see section 5.1 for details on this evaluation framework). This evaluation framework has been used to benchmark the generalization performance of mainstream generative models – Diffusion models [47], GANs [48] and VAEs [49] – against humans in the one-shot drawing setting [30]. Although Diffusion models come closest to human performance, a noticeable gap remained in this study. This article goes further by including representational inductive biases in Latent Diffusion Models, showing these biases significantly narrow the gap with humans.
@@ 3 Datasets
As done in previous work [31, 30, 11], we use the Omniglot [11] and the QuickDraw-FS [30] datasets to compare humans and machines on the one-shot drawing task.
@@ 5 Results
To compare humans and machines in the one-shot drawing task, we first use the originality vs. recognizability framework [31, 30]. This framework leverages $2$ critic networks to evaluate the samples produced during the testing phase. The recognizability is quantified using the classification accuracy of a one-shot classifier [17], while the originality is measured using the average distance between the variations and their corresponding exemplars. This distance is computed in the feature space of a self-supervised model [14]. Note that the originality is normalized across all tested models to range between 00 and $1$. Here, we use the same originality vs. recognizability framework setting as that used in Boutin et al. [30]. Importantly, the originality vs. recognizability plots should be interpreted based on how close the models are to the human data point (grey star in Fig. 2), rather than focusing solely on their individual originality or recognizability scores. In simple terms, a model that effectively mimics human drawings should fall near the human data point.
In Fig. 2, we evaluate how increasing the regularization weights (i.e. the $\beta$ in Eq. 1) for each regularizer affects the similarity of LDM samples to human drawings. To do so, we report the originality and the recognizability values for LDM samples trained with different $\beta$ values (see data points in Fig. 2). We use a parametric fit (least curve fitting methods [54]) to illustrate how increasing $\beta$ affects these scores (see A.5 for more details on the parametric fit computations). We observe a similar concave shape for all curves. As $\beta$ starts increasing, the recognizability improves while the originality decreases (except for VQ regularizer). Beyond a certain $\beta$ value, the recognizability declines, and the originality increases. In particular, the maximum recognizability values for KL and VQ (obtained with $\beta_{KL}=10^{-5}$ and $\beta_{VQ}=5$) match those of a diffusion model trained in the pixel space and barely exceed those of a non-regularized LDM (see Fig. 2a). Increasing the weight of the prototype-based regularizer substantially reduces the distance to human compared to the classification regularizer (the minimal distance to human is $0.04$ for $\beta_{PR}=5\cdot 10^{2}$ vs. $0.15$ for $\beta_{CL}=5$, see Fig. 2b). Among the contrastive regularizers, Barlow regularization significantly reduces the distance to human compared to the SimCLR one (the minimal distance to human is $0.08$ with $\beta_{BAR}=30$ vs. $0.12$ with $\beta_{SimCLR}=10^{-2}$, see Fig. 2c). A visual inspection of the samples tends to corroborate these results (see Fig. 3 and A.7 for more samples). We observe similar trends for all tested regularizers on the Omniglot dataset (see A.6).
Refer to captionExemplars No reg. Proto.Classif.BarlowSimCLRKLVQ
Overall, our findings indicate that while standard regularizers (KL and VQ) slightly enhance the samples’ human likeness, they are not as effective as the other regularizers we tested. Within supervised regularizers, the prototype-based loss proves more beneficial than the classification regularizer. Furthermore, our analysis of contrastive regularizers reveals that reducing feature redundancy between dissimilar samples (as in Barlow regularizer) is more effective at achieving human-level performance than minimizing similarity between samples’ representations (SimCLR).
@@ 6 Conclusion
  • The Barlow objective is the second-best regularizer in the human/LDMs alignment ranking.
@@ 7 Discussion
At first glance, it might be surprising that the self-supervised Barlow regularizer outperforms the supervised classification regularizer. Although the classification loss has access to more external information during training, it may overfit the features of the training categories. This overfitting becomes problematic in few-shot settings, where the training and testing categories differ. Conversely, the prototype-based regularizer, which develops a metric space rather than directly mapping features to labels, does not encounter this problem, suggesting features that are more transferable to the (unseen) testing categories (see  [19, 17] for discussion on this topic). We refer the reader to A.8.6 for a discussion on the limitations of our work.
Figure 5: Effect of combining the prototype-based and Barlow regularizers. The comination corresponds to $\mathcal{L}_{reg}=\beta_{PR}\mathcal{L}_{PR}+\beta_{BAR}\mathcal{L}_{BAR}$, with $\beta_{BAR}=50$ and $\beta_{PR}$ ranging from $0.1$ to $500$. The rest of the caption is similar to Fig. 2.
While this article has considered each regularization independently, there’s potential to combine them, particularly when they offer different representational inductive biases. For example, combining the feature decorrelation of the Barlow objective with the clustering of the prototype-based regularizer could be beneficial. To verify this hypothesis, we have run a (preliminary) experiment in which we leverage this specific combination of regularizers (see Fig. 5). The resulting curve is the closest we have achieved to matching human performance. Interestingly, this hybrid approach enables the LDM to blend the higher originality of the Barlow regularization with the enhanced recognizability from the prototype-based objectives. This finding suggests that smarter combinations of regularization methods could eventually close the performance gap with humans.

### v2

@@ 2 Related work
Human-machine comparison in one-shot drawing tasks: Cognitive scientists have developed various methods to compare the generalization abilities of machines and brains on drawing tasks. Lake et al. [44] introduced the Omniglot challenge in which both humans and machines are tasked with drawing symbols from categories represented by a single exemplar (see [45] for a review on the challenge). The authors evaluated the drawings’ recognizability in a visual Turing test where humans (or classifiers) had to distinguish between human-drawn and machine-generated symbols [11]. Additional metrics, including classification uncertainty and semantic similarity, were also used to compare drawings produced by humans and machines under different time constraints [46, 8]. While these evaluation frameworks provide useful insights into a sample’s recognizability, they do not measure how the diversity of model-generated samples compares to those created by humans. The “originality vs. recognizability” framework [31] mitigates this issue by adding the originality metric. An originality score quantifies the similarity between the original exemplar and its corresponding variations (see section 5.1 for details on this evaluation framework). This evaluation framework has been used to benchmark the generalization performance of mainstream generative models – Diffusion models [47], GANs [48] and VAEs [49] – against humans in the one-shot drawing setting [30]. Although Diffusion models come closest to human performance, a noticeable gap remained in this study. In this article, we use the “originality vs. recognizablility” framework from Boutin et al. [31] to evaluate representational inductive biases in Latent Diffusion Models. In particular, we demonstrate that effective biases in one-shot classification tasks also prove efficient in the one-shot drawing task.
@@ 3 Datasets
As done in previous work [31, 30, 11], we use the Omniglot [11] and the QuickDraw-FS [30] datasets to compare humans and machines on the one-shot drawing task. These datasets, made of handwritten symbols or drawings, offer a fair environment for comparing the generation abilities of humans and machines [11, 46, 31, 30]. It is important to note that natural images generation is a task beyond human capability, making it unsuitable for a fair comparison between humans and machines.
@@ 5 Results
To compare humans and machines in the one-shot drawing task, we first use the originality vs. recognizability framework [31, 30]. This framework leverages $2$ critic networks to evaluate the samples produced during the testing phase. The recognizability is quantified using the classification accuracy of a one-shot classifier [17], while the originality is measured using the average distance between the variations and their corresponding exemplars. This distance is computed in the feature space of a self-supervised model [14]. Importantly, both human-drawn and machine-generated samples are evaluated using the same $2$ critic networks. This ensures that any potential biases in the critic networks are minimized, leading to a more balanced comparative analysis. Note that the originality is normalized across all tested models to range between 00 and $1$. Here, we use the same originality vs. recognizability framework setting as that used in Boutin et al. [30]. Importantly, the originality vs. recognizability plots should be interpreted based on how close the models are to the human data point (grey star in Fig. 3), rather than focusing solely on their individual originality or recognizability scores. In simple terms, a model that effectively mimics human drawings should fall near the human data point. Note also that there is an inherent trade-off between originality and recognizability: while recognizability assesses how likely the data point falls within the classifier decision boundary, originality measures how ’diffuse’ the sample distribution is. Therefore a very original agent (producing highly diverse samples) will tend to have a low recognizability as the samples are likely to fall outside of the classifier decision boundary.
In Fig. 3, we first evaluate how increasing the regularization weights (i.e. the $\beta$ in Eq. 1) for each
Refer to captionExemplars No reg. Proto.Classif.BarlowSimCLRKLVQ
regularizer (taken separately) affects the similarity of LDM samples to human drawings. To do so, we report the originality and the recognizability values for LDM samples trained with different $\beta$ values (see data points in Fig. 3). We use a parametric fit (least curve fitting methods [54]) to illustrate how increasing $\beta$ affects these scores (see A.5 for more details on the parametric fit computations). We observe a similar concave shape for all curves. As $\beta$ starts increasing, the recognizability improves while the originality decreases (except for VQ regularizer). Beyond a certain $\beta$ value, the recognizability declines, and the originality increases. In particular, the maximum recognizability values for KL and VQ (obtained with $\beta_{KL}=10^{-5}$ and $\beta_{VQ}=5$) match those of a diffusion model trained in the pixel space and barely exceed those of a non-regularized LDM (see Fig. 3a). Increasing the weight of the prototype-based regularizer substantially reduces the distance to human compared to the classification regularizer (the minimal distance to human is $0.04$ for $\beta_{PR}=5\cdot 10^{2}$ vs. $0.15$ for $\beta_{CL}=5$, see Fig. 3b). Among the contrastive regularizers, Barlow regularization significantly reduces the distance to human compared to the SimCLR one (the minimal distance to human is $0.08$ with $\beta_{BAR}=30$ vs. $0.12$ with $\beta_{SimCLR}=10^{-2}$, see Fig. 3c). A visual inspection of the samples tends to corroborate these results (see Fig. 2 and A.7 for more samples). We observe similar trends for all tested regularizers on the Omniglot dataset (see A.6).
@@
Overall, our findings indicate that not all regularizers are created equal. For supervised regularizers (see Fig. 3b), the prototype-based regularizer generates more recognizable samples compared to the classification regularizer. This is expected since the classifier focuses on separating categories in the training set, which may not be ideal for unseen categories in the one-shot setting [19, 17]. In contrast, the prototype-based regularizer clusters samples near their prototypes, leading to less overfitting and better transferability, which is valuable for few-shot tasks [55]. Our experiments confirm that the prototype-based regularizer generalizes better for one-shot drawing. In Fig. 3c, the Barlow regularization outperforms the SimCLR regularizer in recognizability, likely due to Barlow’s effective feature disentangling [15]. These features transfer well to new datasets, making Barlow more suitable for the one-shot drawing task. Overall, our results demonstrate that effective representational inductive biases in few-shot learning also enhance performance in one-shot drawing.
We now study the effect of the regularizers when they are used in combination. In particular, we have systematically explored the following combinations of regularizers Barlow + Prototype (Fig. 4a), SimCLR. + Prototype (Fig. 4b), KL + Prototype (Fig. 4c), VQ + Prototype (Fig. 4d). We observe that the Barlow + Prototype and the KL + Prototype combinations produced the most human-like samples. Those regularizer’s combinations are particularly as in both cases the combined recognizability is significantly higher compared to using each regularizer alone. This suggests that clustering samples around their prototypes (using the Prototype regularizer) within a disentangled space (achieved via the KL or Barlow regularizer) enhances generalization. In contrast, the VQ + Prototype and the SimCLR + Prototype combinations show little to no improvements.
Refer to captiona) b)c)d)
Figure 4: Combined effect of the regularization weights on the originality vs recognizability framework (QuickDraw-FS dataset). Each data point represents an LDM trained with a combination of $2$ different regularizers. All combinations include the prototype-based regularizers. The curves represent the parametric fits, oriented in the direction of an increase of $\beta$. a): Barlow and prototype-based regularizers applied either separately (plain lines) or in combination (dashed-line). When applied in combinations, only the weight of the prototype-based regularizer is modified (with $\beta=30$ for Barlow). b): SimCLR and prototype-based regularizers. When applied in combinations, only the weight of the prototype-based regularizer is modified, the SimCLR is set to $\beta=1$. c): KL and prototype-based regularizers. When applied in combinations, only the weight of the prototype-based regularizer is modified, the KL is set to $\beta=1e-3$. d): VQ and prototype-based regularizers. When applied in combinations, only the weight of the prototype-based regularizer is modified, the VQ is set to $\beta=20$. See caption in Fig. 3.
@@ 6 Conclusion
  • In the one-shot drawing tasks, the samples’ human-likeness could be further improved by combining the prototype-based regularizer with either the KL or the Barlow regularizers.
@@ 8 Discussion
Aknowledgement
This work was funded by the European Union (ERC, GLoW, 101096017), ANITI (Artificial and Natural Intelligence Toulouse Institute) and the French National Research Agency (ANR-19-PI3A-0004). Additional funding was provided by ONR (N00014-24-1-2026) and NSF (IIS-1912280, IIS-2402875 and EAR-1925481). Computing hardware supported by NIH Office of the Director grant S10OD025181 via the Center for Computation and Visualization (CCV).